write an equation in slope-intercept form of a line containing the point (4,-5) with a slope equal to -6.
Answer by nerdybill(7384)

write an equation in slope-intercept form of a line containing the point (4,-5) with a slope equal to -6.
.
Plug the given values into the "point-slope" form:
y - y1 = m(x - x1)
where
m is the slope
(x1, y1) is the given point
.
y - y1 = m(x - x1)
y - (-5) = -6(x - 4)
y + 5 = -6x + 24
y = -6x + 19 (this is what they're looking for)